OpenStudy (anonymous):

Find the limit (if it exists). Note that f(x)=|_x_| represents the greatest integer function. lim (x-> -5^+) of {-3|_x_|-9}

OpenStudy (anonymous):

f(x) represents the greatest integer function?? i dint get that?

OpenStudy (anonymous):

yeah I'm not sure what it means either :x

OpenStudy (anonymous):

\(\lfloor x\rfloor=\max\left\{n\in \mathbb{Z}~:~n\le x\right\}\) So the graph looks something like this: